# A Bayesian hierarchical mixture cure modelling framework to utilize multiple survival datasets for long-term survivorship estimates: A case study from previously untreated metastatic melanoma

### v1

@@ 4.2 Prior specification
For the latency model, we centred covariates to aid computation and specified vague priors on the log-scale for the coefficients of the OS and PFS rates, $\log(\lambda_{OS})$ and $\log(\lambda_{PFS})$, respectively. In the baseline scenario these were: ${\beta_{PFS,0}^{\lambda}\sim\text{Normal}(-3,0.5),}$ ${\beta_{OS,0}^{\lambda}\sim\text{Normal}(-3,0.5)}$ and for age ${\beta_{PFS,age}^{\lambda}\sim\text{Normal}(0,0.01),}\;$ ${\beta_{OS,age}^{\lambda}\sim\text{Normal}(0,0.01)}$. The prior distributions for the ancillary parameters $\phi_{OS}$ and $\phi_{PFS}$ depended on the particular survival distribution used. For log logistic and Weibull the prior distribution for the shape parameter was $\phi\sim\text{Gamma}(1,1)$, for Gompertz $\phi\sim\text{Gamma}(1,1000)$, and for log Normal the prior distribution for the standard deviation was $\phi\sim\text{Gamma}(1,2)$.
For the incidence model, the global cure fraction was modelled as a fixed effect linear regression with logistic link. Each treatment coefficient prior is independent ${\beta^{\pi}_{k}\sim\text{Normal}(-0.1,0.2)}$. This corresponds to a prior mean cure fraction of just under 0.5, and approximately 10% and 1% chance of exceeding 0.6 and 0.7 respectively. Alternatively, we could have specified the cure fraction on the natural scale between 0 and 1 using a Beta prior distribution $\pi\sim\text{Beta}(a,b)$. This has the advantage of representing the probability directly which could aid elicitation but has the disadvantage of $a$ and $b$ being less interpretable. The Beta distribution parameters can always be obtained via transformation of the mean and standard deviation. Prior predictive plots and further discussion are provided in the supplementary material.
@@ 4.2.1 Cure fraction between-endpoint variance
For the complete data analysis, the random effect variance on the global cure fraction was modelled using a vague (folded) half-Normal [20] with ${\sigma_{k}\sim\text{Normal}(0,2.5^{2})\mathbb{I}(0,)}$, where $\mathbb{I}(0,)$ denotes a distribution truncated at mean 0. This was found to be the best option for this study. As a rule of thumb, when the number of groups is greater than 5 then a minimally informative uniform prior density associated with the standard deviation $\sigma$ is expected to generally work well, or when there is small hierarchical variance then a $\text{Gamma}(2,0.1)$ may be preferred [21].
As discussed, the early phase of a trial is not very informative of the cure fraction if plateauing is yet to occur. In this situation, we can take advantage of the Bayesian hierarchical structure. In the absence of other evidence, it is reasonable to assume that uncertainty about the cure fractions for each end-point can be modelled using a common distribution, which corresponds to a narrow prior on the between-group standard deviation at the global cure fraction level. In order to do this, we modelled the global cure fraction with a Penalised Complexity (PC) prior [22] that penalizes departure from a base model and preferences simplicity.
In practice, the PC prior approach places an exponential distribution on $\sigma_{k}$ in equation (3). The rate parameter for this distribution, $\rho$, is obtained by first specifying a threshold value $\sigma_{0}$ and probability of exceeding that thereshold, $\alpha$, ie, $p(\sigma>\sigma_{0})=\alpha$. These can then be plugged into $\rho=-\log(\alpha)/\sigma_{0}$, to obtain the rate $\rho$ of the exponential distribution prior. For specific values, for $p(\pi_{k}>0.55)=0.005$ then $\sigma_{0}=0.18$ and $\rho=-\log(0.01)/0.18=25$ or $\sigma\sim\mbox{Exp}(25)$. Similarly, for $p(\pi_{k}>0.525)=0.005$ then $\sigma_{0}=0.08$ and $\rho=-\log(0.01)/0.08=55$ or $\sigma\sim\mbox{Exp}(55)$. The latter prior was used for the data-cut at 12 months. We can see how these two prior distributions correspond to the prior uncertainty on the cure fraction scale in Figure 3. The two distributions are similar, but the plot for $\rho=55$ is more peaked around the central value, preferencing the base model even more.
@@ 4.3 Parametric models for the sampling variability
To model the sampling variability in the observed OS and PFS event times we used the exponential, Weibull, Gompertz, log-normal and log-logistic distributions. These are among those recommended by NICE guidelines for Health Technology Assessments (HTA) [1]. In general, providing this list has been interpreted by many in the HTA literature as meaning all of these distributions should be modelled with a given data set. This prescriptive approach does not take into account what may be known a priori about the problem. In reality, a subset of these parametric models will be more plausible for the data and so these are the ones that should be investigated.
@@ 4.4 Model assessment
We evaluated the goodness of fit of the overall models using out-of-sample predictions estimated with the widely applicable information criterion (WAIC) (and leave-one-out cross-validation, LOO, given in the Appendix) [23]. These have various advantages to more common AIC and DIC. WAIC is the difference between the computed log pointwise predictive density and the estimated effective number of parameters. Smaller values are better. It is fully Bayesian and invariant to parameterisation. They are easily obtained using the posterior sample of the Stan output. The set of parameters in the hierarchical model (and their dimensions) are $\mathbf{\beta^{\pi}}$ (3), $\mathbf{\sigma}$ (3), $\mathbf{\beta^{\lambda}_{PFS}}$ (2), $\mathbf{\beta^{\lambda}_{OS}}$ (2), $\mathbf{\phi_{OS}}$ (1), $\mathbf{\phi_{PFS}}$ (1), $\mathbf{\pi_{OS}}$ (3), $\mathbf{\pi_{PFS}}$ (3).
@@
There has not been much work carried-out to date on model assessment methods specific to mixture cure models. For the basic mixture cure model, assessment may focus on the model as a whole or on the incidence and latency submodels separately. The fit of the latency submodel can be examined using a revised Schoenfeld residual [24], and the performance of the incidence submodel can be checked with concordance measures such as AUC [25]. However, these incidence model approaches are best used directly with known cure status and are designed for models which consider individual cure fractions, unlike our model which assumed a single population cure fraction. Of course, we also extend the basic mixture cure model so it is not yet clear how to best use these methods in this situation.
4.5 60 months data only results
@@
In this section, we will explore the mixture cure model fits using the complete 60-months CheckMate 067 trial data. An example of the Bayesian hierarchical model posterior survival curves using the complete data is shown in Figure 4. Equivalent plots for all other distributions and a combined plot are given in the Appendix.
@@
The background survival is linearly decreasing with time and the same between plots, as expected because were are using point estimates from WHO life tables. Visual inspection demonstrate reasonably good model fits. The PFS Kaplan-Meier curves show an initial steep decrease before levelling off. The exponential fits particularly well to the tail which is of particular importance.
All model fits are fairly similar. For example, for the exponential model, the PFS median times are 5, 9 and 12 months, for ipilimumab, nivolumab and combined treatments, respectively. In comparison with the Kaplan-Meier median times, the fitted times are 75%, 30% and 5% larger. That is, the median times converge for large times. Similarly, for the exponential model, the OS median times are 22, 38 and 56 months, for ipilimumab, nivolumab and combined, respectively. The uncured median survival times for the exponential model are 5 and 15 months, for PFS and OS, respectively.
@@
The restricted mean survival time (RMST) is the average time-to-event over a fixed follow-up period and has been proposed as an intuitive alternative summary. For a follow-up time of 60 months and for the exponential model, the RMSTs (95% credible interval, CrI) for the uncured group are 20.6 (18.4, 23.1) and 6.66 (6.09, 7.39) months, for OS and PFS, respectively. For the OS end-point, the RMSTs are 28.7 (26.3, 31.1), 36.5 (34.1, 38.9) and 39.6 (37.1, 42.1) months, for ipilimumab, nivolumab and combined, respectively. For the PFS end-point, the RMSTs are 11.9 (10.0, 14.1), 24.3 (21.1, 27.4) and 27.7 (24.7, 30.8) months, for ipilimumab, nivolumab and combined, respectively. The RMST for the cured population is the same value corresponding to the general population of 58.2 months.
@@
Figure 5 and Figure 6 show cure fraction posterior distribution forest plots for the hierarchical and separate mixture cure models, respectively. We can see that the global cure fraction posterior is relatively wide, which is understandable since there are only two exchangeable points providing limited information (OS and PFS). They approximately range 0.1-0.55, 0.2-0.6, 0.2-0.65 for ipilimumab, nivolumab and combined treatments respectively. However, despite this, the data influence the end-point cure fractions by pulling the central location away from the prior, downwards for ipilimumab and upwards for the combined treatment. For ipilimumab, the PFS curve fraction posterior is close to containing zero within the credible interval. Whereas for nivolumab and the combined treatment arms the mean global cure fraction lies somewhere between the PFS and OS means, the ipilimumab global mean is closer to the OS cure fraction mean. This is because we have explicitly codified in the prior distribution that the global cure fraction is unlikely to be near zero and the PFS cure fraction posterior is close to zero. We would observe the equivalent behaviour if the OS distributions were near to 1. The survival curves for the whole sample plateau to the background mortality survival from when the uncured survival probability reaches zero. Since the background is user-supplied, in this case from the WHO life tables, then there is no additional insight from this time on-wards. The times at which this occurs are relatively early for PFS and the log-Normal distribution. Comparing the hierarchical and separate models there is not a noticeable difference between the two. As mentioned above, there is limited information with only two end-points and a fairly weak prior. There is a small amount of shrinkage such that the posterior distributions for OS and PFS are pulled toward the global mean, so that the PFS cure fractions are larger and the OS cure fractions are smaller in the hierarchical model than the separate model.
[…]
OS distn PFS distn ELPDnormal-$\dagger$ $p_{D}^{\ddagger}$ WAIC
exp exp -5064.82 (80.38) 11.16 (0.46) 10129.63 (160.77)
Gompertz -5065.64 (80.44) 11.8 (0.49) 10131.29 (160.88)
loglogistic -4970.34 (80.72) 11.78 (0.33) 9940.68 (161.45)
lognormal -4730.47 (92.79) 18.04 (0.91) 9460.94 (185.59)
Weibull -5062.99 (81.69) 14.63 (0.8) 10125.97 (163.38)
Gompertz exp -5064.99 (80.41) 11.45 (0.45) 10129.99 (160.82)
Gompertz -5065.61 (80.44) 11.66 (0.48) 10131.22 (160.88)
loglogistic -4970.42 (80.79) 11.68 (0.34) 9940.84 (161.59)
lognormal -4730.92 (92.64) 18.31 (0.95) 9461.84 (185.29)
Weibull -5063.81 (81.74) 15.45 (0.85) 10127.61 (163.48)
loglogistic exp -5059.86 (80.47) 11.29 (0.42) 10119.73 (160.93)
Gompertz -5060.37 (80.4) 11.64 (0.45) 10120.74 (160.8)
loglogistic -4966.22 (80.75) 12.77 (0.35) 9932.44 (161.5)
lognormal -4726.57 (93.17) 19.43 (0.91) 9453.13 (186.34)
Weibull -5058.17 (81.84) 14.8 (0.74) 10116.33 (163.67)
lognormal exp -4972.55 (84.58) 16.35 (0.61) 9945.09 (169.15)
Gompertz -4973.16 (84.69) 16.5 (0.56) 9946.32 (169.38)
loglogistic -4877.87 (85.79) 16.53 (0.5) 9755.74 (171.59)
lognormal -4638.85 (100.27) 23.58 (1) 9277.71 (200.55)
Weibull -4970.22 (86.17) 19.01 (0.89) 9940.44 (172.34)
Weibull exp -5063.26 (80.7) 11.93 (0.46) 10126.53 (161.4)
Gompertz -5064.37 (80.64) 12.67 (0.52) 10128.73 (161.28)
loglogistic -4968.74 (81.03) 12.3 (0.37) 9937.49 (162.07)
lognormal -4729.78 (93.18) 19.54 (0.97) 9459.56 (186.36)
Weibull -5061.91 (82.05) 15.77 (0.85) 10123.81 (164.11)
[…]
OS distn PFS distn ELPDnormal-$\dagger$ $p_{D}^{\ddagger}$ WAIC
exp exp -5066.38 (81.4) 10.46 (0.43) 10132.76 (162.81)
Gompertz -5066.68 (81.46) 10.38 (0.46) 10133.37 (162.92)
loglogistic -4971.73 (81.67) 10.36 (0.28) 9943.46 (163.34)
lognormal -4732.41 (93.18) 18.56 (0.99) 9464.83 (186.36)
Weibull -5064.87 (82.86) 14.39 (0.86) 10129.74 (165.72)
Gompertz exp -5066.72 (81.44) 10.51 (0.45) 10133.43 (162.89)
Gompertz -5066.76 (81.49) 10.19 (0.44) 10133.53 (162.97)
loglogistic -4972.24 (81.7) 10.7 (0.28) 9944.47 (163.4)
lognormal -4732.23 (93.3) 18.01 (0.95) 9464.46 (186.61)
Weibull -5064.46 (82.92) 13.5 (0.78) 10128.91 (165.83)
loglogistic exp -5062.51 (81.89) 10.79 (0.4) 10125.02 (163.78)
Gompertz -5062.64 (82.01) 10.26 (0.41) 10125.27 (164.02)
loglogistic -4968.45 (82.36) 11.63 (0.28) 9936.89 (164.72)
lognormal -4728.25 (94.2) 18.18 (0.92) 9456.51 (188.4)
Weibull -5060.18 (83.31) 13.79 (0.69) 10120.37 (166.62)
lognormal exp -4973.43 (85.26) 14.63 (0.55) 9946.86 (170.53)
Gompertz -4973.48 (85.22) 14.31 (0.51) 9946.96 (170.44)
loglogistic -4879.07 (86.55) 15.07 (0.44) 9758.14 (173.09)
lognormal -4638.43 (100.4) 21.3 (0.96) 9276.85 (200.8)
Weibull -4972.48 (87.07) 19.54 (0.91) 9944.95 (174.13)
Weibull exp -5064.74 (81.52) 11.11 (0.45) 10129.48 (163.04)
Gompertz -5065.16 (81.59) 11.16 (0.44) 10130.32 (163.18)
loglogistic -4970.64 (81.87) 11.68 (0.35) 9941.27 (163.73)
lognormal -4730.93 (93.78) 19.04 (0.96) 9461.85 (187.56)
Weibull -5063.2 (83.09) 14.87 (0.83) 10126.41 (166.17)
Figure 5: Hierarchical mixture cure model posterior cure fraction forest plots with 95% credible intervals for (i) ipilimumab only (ii) nivolumab only and (iii) combination treatment.
Figure 6: Separate mixture cure models posterior cure fraction forest plots with 95% credible intervals for (i) ipilimumab only (ii) nivolumab only and (iii) combination treatment.
4.6 All data-cuts results
@@
We now turn our attention to fitting the mixture cure models to earlier data-cuts in the CheckMate 067 trial data set. For the complete data set up to 60 months we have seen that the separate and hierarchical models produce similar parameter inferences. This is to be expected when the data are relatively mature, there are in effect only two data points at the global level and a weak prior is used for the global variance hyper parameter.
For demonstration purposes, we show results using the penalised complexity prior with exponential distribution rate parameter $\lambda=55$ for the 12 months data cut scenario and retain all other previous priors, including the weak priors for the mean $\pi_{k}$. More data are available at later times and so the emphasis on this prior was relaxed. This approach is in the spirit of information-sharing methods such as power priors, which are ”adaptively robust”, where a power parameter controls the degree of borrowing. For example, for the simple case of normal historical data under a flat prior, the power prior is another normal distribution with a scaled variance according to the sample size and magnitude of the power parameter, determined by the investigator.
Figures 7 and Figure 8 show forest plots of the cure fraction estimates for 12, 30 and 60 months, and for the separate and hierarchical models, respectively. Results are shown for the OS exponential with PFS exponential model and the OS log-Normal with PFS log-Normal model.
@@
We can see that in Figure 7 for the exponential model in the OS survival curves for the separate model there is more uncertainty than in the equivalent plot for the hierarchical model in Figure 8. The log-Normal model OS survival curves for separate and hierarchical models over-estimate the cure fraction at earlier data-cuts and all treatment arms. For later data-cuts, the cure fractions appear to converge from above to the complete data estimates.
@@
Figures 9 shows the survival curves for the data-cut with 12 months of follow-up data for the exponential and for the separate and hierarchical models, respectively. Equivalent Figures for 30 months data-cut are given in the Appendix.
@@
Figure 7: Posterior cure fraction forest plot with 95% credible intervals for study data with cutpoint censoring at 12 months, 30 months and the complete data set using the separate mixture cure models.
Figure 8: Posterior cure fraction forest plot with 95% credible intervals for study data with cutpoint censoring at 12 months, 30 months and the complete data set using the hierarchical mixture cure model.
@@
Table 3 shows the survival estimates at 60 months extrapolated using 12 and 30 months data cut point. Clearly, when the survival curve has converged to the long-term plateau this is equivalent to the cure fraction.
Table 3: Survival probability estimates for separate and hierarchical models at 60 months using 12, 30 and 60 months data cut points. Models with exponential or log-normal distributions for both OS and PFS are given.
Endpoint Model Distribution Cut-point Ipilimumab Nivolumab Combined
OS hier exp 12 0.19 [0.14, 0.28] 0.34 [0.28, 0.39] 0.42 [0.33, 0.48]
OS hier exp 30 0.2 [0.11, 0.28] 0.33 [0.28, 0.38] 0.47 [0.4, 0.55]
OS hier exp 60 0.25 [0.2, 0.29] 0.31 [0.26, 0.36] 0.51 [0.46, 0.56]
OS hier lognormal 12 0.28 [0.23, 0.33] 0.38 [0.33, 0.43] 0.49 [0.43, 0.54]
OS hier lognormal 30 0.23 [0.17, 0.3] 0.33 [0.28, 0.38] 0.49 [0.43, 0.56]
OS hier lognormal 60 0.25 [0.21, 0.3] 0.31 [0.26, 0.35] 0.51 [0.46, 0.56]
OS separate exp 12 0.26 [0.18, 0.34] 0.34 [0.28, 0.4] 0.34 [0.23, 0.44]
OS separate exp 30 0.22 [0.17, 0.28] 0.32 [0.27, 0.38] 0.47 [0.4, 0.52]
OS separate exp 60 0.25 [0.21, 0.3] 0.3 [0.25, 0.35] 0.5 [0.45, 0.55]
OS separate lognormal 12 0.34 [0.26, 0.42] 0.38 [0.32, 0.43] 0.41 [0.33, 0.51]
OS separate lognormal 30 0.27 [0.22, 0.32] 0.33 [0.27, 0.37] 0.48 [0.42, 0.54]
OS separate lognormal 60 0.27 [0.22, 0.32] 0.3 [0.25, 0.35] 0.51 [0.45, 0.56]
PFS hier exp 12 0.1 [0.05, 0.15] 0.39 [0.34, 0.44] 0.38 [0.32, 0.44]
PFS hier exp 30 0.1 [0.07, 0.14] 0.41 [0.34, 0.49] 0.39 [0.33, 0.45]
PFS hier exp 60 0.08 [0.05, 0.12] 0.42 [0.37, 0.48] 0.38 [0.32, 0.44]
PFS hier lognormal 12 0.16 [0.11, 0.2] 0.47 [0.41, 0.51] 0.43 [0.38, 0.48]
PFS hier lognormal 30 0.1 [0.07, 0.14] 0.43 [0.38, 0.49] 0.4 [0.34, 0.45]
PFS hier lognormal 60 0.08 [0.05, 0.12] 0.43 [0.38, 0.48] 0.38 [0.32, 0.43]
PFS separate exp 12 0.1 [0.06, 0.15] 0.35 [0.25, 0.45] 0.4 [0.33, 0.46]
PFS separate exp 30 0.11 [0.08, 0.15] 0.41 [0.34, 0.48] 0.38 [0.32, 0.43]
PFS separate exp 60 0.1 [0.07, 0.14] 0.42 [0.37, 0.48] 0.37 [0.31, 0.42]
PFS separate lognormal 12 0.15 [0.11, 0.19] 0.43 [0.35, 0.52] 0.44 [0.39, 0.49]
PFS separate lognormal 30 0.12 [0.08, 0.15] 0.43 [0.36, 0.5] 0.39 [0.34, 0.45]
PFS separate lognormal 60 0.1 [0.07, 0.13] 0.43 [0.38, 0.48] 0.37 [0.31, 0.42]

### v2

@@ 5.2 Prior specification
For the latency model, we centred covariates to aid computation and specified minimally informative priors on the log-scale for the coefficients of the OS and PFS rates, $\log(\lambda_{OS})$ and $\log(\lambda_{PFS})$, respectively. These were the same for both OS and PFS: ${\beta_{\cdot,0}^{\lambda}\sim\text{Normal}(-3,0.5)}$ and for age $\beta_{\cdot,age}^{\lambda}\sim\text{Normal}(0,0.01)$. These values were picked to be vague about the true rate but the median survival time was expected to have been reached before 30 months. Figure A4 in the Appendix shows realisations of survival curves using samples from the prior for $\beta_{\cdot,0}^{\lambda}$. At 30 months the median survival probability is 0.22 [IQR: 0.01 - 0.58] and at 60 months is 0.05 [IQR: 0 - 0.34]. The prior distributions for the ancillary parameters $\phi_{OS}$ and $\phi_{PFS}$ depended on the particular survival distribution used. For log logistic and Weibull the prior distribution for the shape parameter was $\phi\sim\text{Gamma}(1,1)$, for Gompertz $\phi\sim\text{Gamma}(1,1000)$, and for log Normal the prior distribution for the standard deviation was $\phi\sim\text{Gamma}(1,2)$.
For the incidence model, the global cure fraction was modelled as a fixed effect linear regression with logistic link. Each treatment coefficient prior is independent $\beta^{\pi}_{k}\sim\text{Normal}(-0.1,0.2)$. This corresponds to a prior mean cure fraction of just under 0.5, and approximately 10% and 1% chance of exceeding 0.6 and 0.7, respectively. This is a relatively weak prior (see simulation study) but still informed by expert judgement. Alternatively, we could have specified the cure fraction on the natural scale between 0 and 1 using a Beta prior distribution $\pi\sim\text{Beta}(a,b)$. This has the advantage of representing the uncertainty about the parameter directly which could aid elicitation but has the disadvantage of $a$ and $b$ being less interpretable. The Beta distribution parameters can always be obtained via transformation of the mean and standard deviation of the Normal distribution. Prior predictive distribution plots and further discussion are provided in the Appendix.
@@ 5.2.1 Cure fraction between-endpoint variance
For the complete data analysis, the random effect variance on the global cure fraction was modelled using a minimally informative (folded) half-Normal 22 with ${\sigma_{k}\sim\text{Normal}(0,2.5^{2})\mathbb{I}(0,)}$, where $\mathbb{I}(0,)$ denotes a distribution truncated at mean 0. We considered several alternatives following the recommendations in Gelman (2006) and this was found to be the best option for this study because it balances vagueness and support at zero. As a rule of thumb, when the number of groups is greater than 5 then a minimally informative uniform prior density associated with the standard deviation $\sigma$ is expected to generally work well, or when there is small hierarchical variance then a $\text{Gamma}(2,0.1)$ may be preferred 23.
As discussed, the early phase of a trial is not very informative of the cure fraction if plateauing is yet to occur. In this situation, we can take advantage of the Bayesian hierarchical structure. In the absence of other evidence, it is reasonable to assume that uncertainty about the cure fractions for each end-point can be modelled using a common distribution, which corresponds to a narrow prior on the between-group standard deviation at the global cure fraction level. In order to do this, we modelled the global cure fraction with a Penalised Complexity (PC) prior 24 that penalizes departure from a base model and preferences simplicity. That is, the base model as the pooled data model is the most parsimonious model.
In practice, the PC prior approach places an exponential distribution on $\sigma_{k}$ in equation (3). The rate parameter for this distribution, $\rho$, is obtained by first specifying a threshold value $\sigma_{0}$ and probability of exceeding that threshold, $\alpha$, ie, $p(\sigma>\sigma_{0})=\alpha$. These can then be plugged into $\rho=-\log(\alpha)/\sigma_{0}$, to obtain the rate $\rho$ of the exponential distribution prior. For specific values, for $p(\pi_{k}>0.55)=0.005$ then $\sigma_{0}=0.18$ and $\rho=-\log(0.01)/0.18=25$ or $\sigma\sim\mbox{Exp}(25)$. Similarly, for $p(\pi_{k}>0.525)=0.005$ then $\sigma_{0}=0.08$ and $\rho=-\log(0.01)/0.08=55$ or $\sigma\sim\mbox{Exp}(55)$. The latter prior was used for the data-cut at 12 months. When more data are available at later times the use of this prior was replaced by the truncated half-Normal as discussed elsewhere. This approach is in the spirit of information-sharing methods such as power priors, which are "adaptively robust", where a power parameter controls the degree of borrowing. For example, for the simple case of normal historical data under a flat prior, the power prior is another normal distribution with a scaled variance according to the sample size and magnitude of the power parameter, determined by the investigator. We can see how the two prior distributions correspond to the prior uncertainty on the cure fraction scale in Figure 3. The two distributions are similar, but the plot for $\rho=55$ is more peaked around the central value, preferencing the base model even more.
@@ 5.3 Parametric models for the sampling variability
To model the sampling variability in the observed OS and PFS event times we used the exponential, Weibull, Gompertz, log-normal and log-logistic distributions. These are among those recommended by NICE guidelines for Health Technology Assessments (HTA) 1. In general, providing this list has been interpreted by many in the HTA literature as meaning all of these distributions should be modelled with a given data set. This prescriptive approach does not take into account what may be known a priori about the problem. In reality, a subset of these parametric models will be more plausible for the data and so these are the ones that should be investigated.
@@ 5.4 Model assessment
We evaluated the goodness of fit of the overall models using out-of-sample predictions estimated with the widely applicable information criterion (WAIC) (and leave-one-out cross-validation, LOO, given in the Appendix) 25. These have various advantages to more common AIC and DIC. WAIC is the difference between the computed log pointwise predictive density and the estimated effective number of parameters. Smaller values are better. It is fully Bayesian and invariant to parameterisation. They are easily obtained using the posterior sample of the Stan output.
@@
There has not been much work carried-out to date on model assessment methods specific to mixture cure models. For the basic mixture cure model, assessment may focus on the model as a whole or on the incidence and latency submodels separately. The fit of the latency submodel can be examined using a revised Schoenfeld residual 26, and the performance of the incidence submodel can be checked with concordance measures such as AUC 27. However, these incidence model approaches are best used directly with known cure status and are designed for models which consider individual cure fractions, unlike our model which assumed a single population cure fraction. Of course, we also extend the basic mixture cure model so it is not yet clear how to best use these methods in this situation.
5.5 Results using the complete data set at 60 months
@@
In this section, we will explore the mixture cure model fits using the complete 60-months CheckMate 067 trial data. An example of the Bayesian hierarchical model posterior survival curves using the complete data and exponential distributions for both OS and PFS is shown in Figure 4. We show this because, even though the exponential is the simplest of the NICE recommend distributions, from visual inspection it fits fairly well. Equivalent plots for all other distributions and a combined plot are given in the Appendix.
@@
The background survival appears to decrease at a small, steady rate and is the same between plots, as expected because the point estimates from WHO life tables were used directly. Visual inspection demonstrates reasonably good model fits. The PFS Kaplan-Meier curves show an initial steep decrease before levelling off. The exponential fits particularly well to the tail which is of particular importance because this is usually where extrapolations have the worst fit.
@@
For a follow-up time of 60 months and for the exponential model, the RMSTs (95% credible interval, CrI) for the uncured group are 20.6 (18.4, 23.1) and 6.66 (6.09, 7.39) months, for OS and PFS, respectively. For the OS end-point, the RMSTs are 28.7 (26.3, 31.1), 36.5 (34.1, 38.9) and 39.6 (37.1, 42.1) months, for ipilimumab, nivolumab and combined, respectively. For the PFS end-point, the RMSTs are 11.9 (10.0, 14.1), 24.3 (21.1, 27.4) and 27.7 (24.7, 30.8) months, for ipilimumab, nivolumab and combined, respectively. The RMST for the cured population is the same value corresponding to the general population of 58.2 months.
@@
The global cure fraction posterior is relatively wide, which is understandable since there are only two exchangeable points providing limited information (OS and PFS). They range 0.1-0.55, 0.2-0.6, 0.2-0.65 for ipilimumab, nivolumab and combined treatments respectively. However, despite this, the data influence the end-point cure fractions by pulling the central location away from the prior, downwards for ipilimumab and upwards for the combined treatment. For ipilimumab, the PFS curve fraction posterior is close to containing zero within the credible interval. Whereas for nivolumab and the combined treatment arms the mean global cure fraction lies somewhere between the PFS and OS means, the ipilimumab global mean is closer to the OS cure fraction mean. This is because we have explicitly codified in the prior distribution that the global cure fraction is unlikely to be near zero and the PFS cure fraction posterior is close to zero. We would observe the equivalent behaviour if the OS distributions were near to 1. The survival curves for the whole sample plateau to the background mortality survival from when the uncured survival probability reaches zero. Since the background is user-supplied, in this case from the WHO life tables, then there is no additional insight from this time on-wards. The times at which this occurs are relatively early for PFS and the log-Normal distribution. Comparing the hierarchical and separate models there is not a noticeable difference between the two. As mentioned above, there is limited information with only two end-points and a fairly weak prior. There is a small amount of shrinkage such that the posterior distributions for OS and PFS are pulled toward the global mean, so that the PFS cure fractions are larger and the OS cure fractions are smaller in the hierarchical model than the separate model.
5.6 Results using all data-cuts
@@
We now turn our attention to fitting the mixture cure models to earlier data-cuts in the CheckMate 067 trial data set. For the complete data set up to 60 months we have seen that the separate and hierarchical models produce similar parameter inferences. This is to be expected when the data are relatively mature, there are in effect only two data points at the global cure fraction level and a weak prior is used for the global variance hyper parameter. For demonstration purposes, we show results using the penalised complexity prior with exponential distribution rate parameter $\lambda=55$ for the 12 months data cut scenario and retain all other previous priors, including the weak priors for the mean $\pi_{k}$. Recall that for later times, when more data are available, the alternative variance prior was adopted.
@@
Figures 5a and Figure 5b show forest plots of the cure fraction estimates for 12, 30 and 60 months, and for the separate and hierarchical models, respectively. Results are shown for the OS exponential with PFS exponential model and the OS log-Normal with PFS log-Normal model. We can see that in Figure 5a for the exponential model in the OS survival curves for the separate model there is more uncertainty than in the equivalent plot for the hierarchical model in Figure 5b. The log-Normal model OS survival curves for separate and hierarchical models over-estimate the cure fraction at earlier data-cuts and all treatment arms. For later data-cuts, the cure fractions appear to converge from above to the complete data estimates.
@@
Figure 6 shows the survival curves for the data-cut with 12 months of follow-up data for the exponential and for the separate and hierarchical models, respectively. Figures for 30 months data-cut and table of the survival estimates at 60 months extrapolated using 12 and 30 months data cut point are given in the Appendix.
@@
(a) Separate Mixture Cure Models
(b) Hierarchical Mixture Cure Model
Figure 5: Posterior cure fraction forest plots with 95% credible intervals for study data with cutpoint censoring at 12 months, 30 months, and the complete dataset.